OpenStudy (anonymous):

Set up integrals for both orders of integration and use the more convenient order to evaluate the integral over the region R. R: region bounded by y=0,y=x(squarert), x = 4 y/(1+x^2) dA

OpenStudy (anonymous):

I'm going to retype the eqution

OpenStudy (anonymous):

Why dont you set them up >:(

OpenStudy (anonymous):

Lol jk

OpenStudy (anonymous):

\[\int\limits_{R}^{}\int\limits_{}^{}y/(y/(1+x^2) dA\] These are the given limits: \[y=0,y=\sqrt{x},x=4\]

OpenStudy (anonymous):

opps the numerator should be 1 not y

OpenStudy (anonymous):

\[\int\limits_{R}^{}\int\limits_{}^{}y/1+x^2dA\]

OpenStudy (anonymous):

Since I'm not given the second x limit, I'm assuming I'll have to graph y=\[\sqrt{x}\] and x=4